# Polarizable Water Model with Ab Initio Neural Network Dynamic Charges and Spontaneous Charge Transfer

### v1

@@ 2.2 Water model with ChargeNN charges
The potential energy functional of the 3-site flexible water model incorporating ChargeNN charges and charge transfer contributions is expressed as
@@
$\begin{split}U^{\mathrm{intra}}{}&=\sum_{I}\left[\sum_{i}\frac{k_{b}}{2}(r_{\text{O}_{I}\textbf{H}_{iI}}-r^{0}_{\text{OH}})^{2}+\frac{k_{a}}{2}(\theta_{I}-\theta^{0})^{2}\right]\\ U^{\mathrm{inter}}{}&=\sum_{i<j}\varepsilon_{ij}\left[7\left(\frac{\sigma_{ij}}{r_{ij}}\right)^{9}-5\left(\frac{\sigma_{ij}}{r_{ij}}\right)^{6}\right]\\ {}&+\sum_{i<j}k_{e}\frac{q_{i}q_{j}}{r_{ij}}-\sum_{I<J}k_{c}\frac{\delta q_{IJ}^{2}}{\delta q_{IJ}^{2}+3(\delta q^{0})^{2}}.\end{split}$ (5)
@@
We observed that atomic charges exhibit significant sensitivity to stretching bonds and bending angles, necessitating the addition of flexibility in the intramolecular forces. Moreover, this flexibility has been reported to improve predicted water properties, and also permits recovering intramolecular vibrational spectra.58, 73 Consequently, we adopted harmonic bonds and angles for better description of the dynamical behavior of the water molecules. The equilibrium bond length $r^{0}_{\text{OH}}$ and angle $\theta^{0}$ were tuned to ensure that the average values in the liquid phase align with experimental measurements at 298.15 K and 1 atm. The coefficients of bond length ($k_{b}$) and angle ($k_{a}$) were adjusted to produce accurate positions of the bond stretching and bending peaks in the IR spectrum.
@@
The electrostatic energy is calculated using ChargeNN charges. To facilitate simulations under periodic boundary conditions (PBC), Ewald summation is employed, utilizing the minimum image convention to construct the local environment for charge generation. It is important to note that all atomic charges within a cluster or unit cell must be subtracted by a small constant of approximately $\pm$0.0001 e to ensure a zero net charge of the whole system. There have been attempts to include charge transfer interactions with empirical charge transfers.118, 119, 120, 121 For our model that uses CT obtained quantum mechanically, the charge transfer term formulates a smooth ratio function of squared CT, where $\delta q^{0}$ denotes the reference CT that yields the strongest force. The formula offers several merits, for instance, it is symmetric and the CT energy converges to zero in the absence of CT, preventing abrupt changes in potential energy around the CT cutoff. Additionally, the formula effectively captures the rapid intensification of CT interactions when initial charges are small, as well as the gradual attenuation due to the electron migration from equilibrium. Moreover, since the inter-water CT occurs away from the nuclear sites, the CT term naturally accounts for out-of-plane polarization, resulting in more accurate water conformations.
@@
As the inclusion of the CT term strengthens the short-range interactions, the traditional 12-6 Lennard-Jones (LJ) potential for Van der Waals (VDW) forces yielding a steep repulsive wall constrains inter-water distances to a compact range, which leads to an excessively high first peak in the oxygen-oxygen radial distribution function (RDF). Instead, we employ a 9-6 functional form with minor revisions to the original LJ formula, retaining the 12-6 LJ parameters $\varepsilon$ and $\sigma$ for compatibility with the existing force fields. In the process of the parameter tuning, the VDW equation was written as $U_{ij}^{\mathrm{vdw}}=A_{ij}/r^{9}_{ij}-B_{ij}/r^{6}_{ij}$ for convenience. The intermolecular parameters $k_{c}$, $k_{d}$, $A_{ij}$ and $B_{ij}$ were optimized for the accurate heat of vaporization, density and radial distribution functions (RDF) for liquid water at the ambient conditions. All parameters can be found in Table S1.
@@ 2.3 Implementation
Figure 1 illustrates the implementation of the ChargeNN water model for molecular dynamics (MD) simulations. In each MD step, atomic coordinates adjusted by the MD solver are utilized to compute the ML features interaction-classified functions $\mathbf{F}$ and their analytical derivatives with respect to nuclear positions $\mathrm{d}\mathbf{F}/\mathrm{d}\mathbf{R}$. The ML features are then fed into the pre-trained neural networks for the predictions of atomic charges and charge transfers. By employing auto-differentiation through back-propagation, the gradients of the charges concerning the features $\mathrm{d}q/\mathrm{d}\mathbf{F}$ can be conveniently evaluated. Subsequently, the predicted atomic charges and charge transfers are used to determine the potential energy as described in eq. 5, where total gradients are obtained using the chain rule:
@@
The algorithm has been parallelized based on Message Passing Interface (MPI) standard of version 3. The shared memory windows built in MPI-3’s remote memory access module enable rapid data communications between processes, with significantly lower latency by avoiding excessive memory copies and synchronizations inherent in traditional point-to-point communications. The algorithm of the ChargeNN water model can be readily adapted to CUDA or OPENCL for massive parallelization on graphics processing units (GPUs) in our future implementation.
@@ 3.1 Performance of the charge model
The ChargeNN model demonstrates exceptional data efficiency, requiring only 8000 training (H2O)25 clusters to achieve converged mean absolute errors (MAEs) for the remaining testing clusters, as shown in Figure S3. According to Figure 2, when trained on 8000 (H2O)25, the charge model yields testing MAEs of 0.0025 e for hydrogen and 0.0039 e for oxygen within the same system. These values are markedly lower than the respective MAEs of 0.0036 e and 0.0045 e for (H2O)23 predicted by a recent KRR water charge model trained with the same number of data points32. In addition, our model shows exceptionally high prediction reliability with the coefficients of determination ($R^{2}$) of 0.96 for hydrogen and 0.93 for oxygen, outperforming the 0.88 reported for the KRR model32. Notably, the ChargeNN model considerably improves the transferability, only with a slow deterioration in accuracy as the water cluster size increases. As demonstrated in Figure 2, the models trained with (H2O)25 make accurate prediction of the charges for much larger (H2O)190, yielding MAEs of 0.0028 e for hydrogen and 0.0042 e for oxygen, both with $R^{2}$ values above 0.9. Figure S4 shows that predicting molecular charges by summing the ChargeNN atomic charges is less satisfactory, with $R^{2}>0.92$ and marginally larger MAEs of $\sim$0.007 e due to the accumulative errors, which can be improved with an additional model for molecular charges in future. For assessing ChargeNN’s performance in predicting charge transfer, we prepared charge transfers of 40000 water pairs extracted from the MD trajectory of (H2O)25. The ChargeNN trained on 20000 water pairs achieves an MAE of only 0.001 e and $R^{2}$ of 1.00 for the remaining testing pairs.
@@ 3.2 Performance of the water model
aThe origin was set to be the center of the nuclear charge of each molecule. bMolecules were wrapped to the unit cell. cThe gas-phase properties, interaction energy and melting temperature for SPC/FW were computed in this work, while other data were from ref.58. dSWM4-NDP results from ref.72, 73. eExperimental results from ref.122, 123, 124, 125, 126, 127.
@@
To assess the performance of the ChargeNN water model, we calculated various properties of water in gas, liquid, and solid states. Computational details are provided in the SI. We compare our results with those of SPC/FW58, which is the fixed charge counterpart of the 3-site ChargeNN model, and those with a widely used polarizable model SWM4-NDP72. As shown in Table 1, the ChargeNN’s predictions of water monomer and dimer properties agree well with experimental results, demonstrating its capability to accurately simulate gas-phase water clusters. Figure 3a illustrates that the angle $\theta_{A}$ of the water dimer obtained with ChargeNN (65) closely matches the experimental value of 58, in contrast to the unreasonable 22 by SPC/FW, highlighting the significant impact of polarization and charge transfer on the equilibrium geometries.
@@
Table 1 also demonstrates that the ChargeNN water model accurately reproduces bulk water properties at room temperature and standard pressure, such as average bond length and angle, density, interaction energy, vaporization heat, and diffusion constants. As shown in Figure 3b, ChargeNN produces an RDF for liquid oxygen-oxygen distances in good agreement with the MP2128 and early experimental data129. Nevertheless, the latest x-ray diffraction experiment130 has measured a less pronounced first peak of approximately 2.57 than the predicted one of 2.98, indicating the need for further parameter refinement to mitigate the “over-structured” hydrogen-bond network. The predicted infrared spectra in Figure 3c were obtained through a Fourier transformation of the autocorrelation function of the total dipole moments, in which the one produced by ChargeNN matches the experimental spectrum closely in both band positions and intensities. Notably, ChargeNN successfully captures the hydrogen-bond stretching peak at $\sim$200 cm-1, which is absent in the spectra by SPC/FW and SWM4-NDP. Our results and a previous study32 indicate that accurately describing hydrogen-bond stretching requires precise polarization and intermolecular CT. Additionally, ChargeNN retrieves the bending-libration combination band at $\sim$2150 cm-1, which is missing in the SPC/FW spectrum, further underscoring the importance of explicit polarization in predicting vibrational properties. Nevertheless, we observe that the O-H stretching band at around 3400 cm-1 is considerably narrower than the reference, which may be improved by replacing the harmonic intramolecular forces with anharmonic forces63 or reactive force fields131, 132.
@@
Figure 3: Assessment of the ChargeNN water model. (a) Geometries of water dimers optimized by SPC/FW and ChargeNN, compared to the experimental structure123. (b) Liquid radial distribution functions of the oxygen-oxygen distance obtained with ChargeNN, MP2128 and experiments129, 130 under ambient conditions. (c) Liquid infrared spectra generated by SPC/FW, SWM4-NDP32, ChargeNN and experiment133 under ambient conditions. (d) Temperature dependent densities from SPC/FW134, SWM4-NDP74, ChargeNN and experiment135 at 1 atm.
@@
We computed the molecular dipole moments by setting the origin as the nuclear charge center of each partially charged water molecule, following the reported protocol115, 32. While ChargeNN accurately reproduces dipole moments for water monomer and dimer, as well as the liquid IR spectrum, it yields an average molecular dipole moment of approximately 1.76 Debye, which is notably lower than the experimental estimation of 2.9 Debye. Given that MP2 has been shown to accurately predict water dipole moments115, we validated the accuracy of dipole moments obtained from CM5 charges by comparing them with the MP2 reference for Cambridge water clusters comprising 2 to 20 molecules136, using the MBE(3)-OSV-MP2 relaxed density matrix with the aug-cc-pVTZ basis set. As presented in Figure S7, the CM5 dipole moments are in excellent agreement with the MP2 reference for all tested water clusters, with only minor deviations. Thus, the discrepancy between the predicted average molecular dipole moment and the experimental value may be attributed to the origin-dependent nature of the non-neutral molecules. Additionally, our water model underestimates the dielectric constant obtained from the total liquid dipole moments, which is largely due to the non-uniqueness of the dipole moment of an extended system according to the modern theory of polarization137.
@@
The accurate characterization across a range of temperatures is important to the performance of a water model. Figure 3d demonstrates that ChargeNN successfully predicts the temperature-dependent density of liquid water, with the temperature of maximum density around 280 K, very close to the experimental value of 277 K. In contrast, this prediction remains a challenge for the fixed charge model SPC/FW and other well-tuned 3-site models57, emphasizing the substantial improvement resulting from ChargeNN polarization. The polarizable SWM4-NDP also falls short in predicting the density variation with temperature. Additionally, Figure S6 presents that the computed temperature-dependence of vaporization enthalpy by ChargeNN agrees well with the reference. For the simulation of ice, we calculated the orientational tetrahedral order parameters of the equilibrium configurations across various temperatures, starting from a perfect hexagonal ($I_{h}$) ice structure. The ChargeNN yields a melting temperature of 288 K, close to the experimental value of 273 K. While this result is significantly better than SPC/FW at 190 K and SWM4-NDP at 185 K73, there is still room to fine-tune the parameters to achieve ice properties comparable to TIP4P/Ice54.

### v3

@@ 2.3 Water model with ChargeNN charges
The potential energy functional of the 3-site flexible water model incorporating ChargeNN charges and charge transfer contributions is expressed as a sum of the intramolecular energy ($U^{\mathrm{intra}}$) and the intermolecular energy ($U^{\mathrm{inter}}$),
@@
$\begin{split}U^{\mathrm{intra}}{}&=\sum_{I}\left[\sum_{i}\frac{k_{b}}{2}(r_{\text{O}_{I}\text{H}_{i_{I}}}-r^{0}_{\text{OH}})^{2}+\frac{k_{a}}{2}(\theta_{I}-\theta^{0})^{2}\right],\\ U^{\mathrm{inter}}{}&=\sum_{i<j}\varepsilon_{ij}\left[7\left(\frac{\sigma_{ij}}{r_{ij}}\right)^{9}-5\left(\frac{\sigma_{ij}}{r_{ij}}\right)^{6}\right]\\ {}&+\sum_{i<j}k_{e}\frac{q_{i}q_{j}}{r_{ij}}-\sum_{I<J}k_{c}\frac{\delta q_{IJ}^{2}}{\delta q_{IJ}^{2}+3(\delta q^{0})^{2}}.\end{split}$ (9)
In the formula for intramolecular energy, $k_{b}$ and $k_{a}$ represent the coefficients of bond stretching and bending, respectively. $r_{\text{O}_{I}\text{H}_{i_{I}}}$ denotes the O-H bond length, while $\theta_{I}$ refers to the H-O-H angle of the molecule $I$, with $r^{0}_{\text{OH}}$ and $\theta^{0}$ being the equilibrium values. For the intermolecular energy, $\epsilon_{ij}$ and $\sigma_{ij}$ are standard 12-6 Lennard-Jones (LJ) parameters. $k_{e}$ and $k_{c}$ pertain to coulomb constant and charge-transfer constant, respectively. $q_{i}$ denotes the charge of atom $i$. $\delta q_{IJ}$ represents the charge transferred between molecules $I$ and $J$, with the parameter $\delta q^{0}$ as the reference CT that yields the strongest CT force.
@@
We observed that atomic charges exhibit significant sensitivity to stretching bonds and bending angles, necessitating the addition of flexibility in the intramolecular forces. Moreover, this flexibility has been reported to improve predicted water properties, and also permits recovering intramolecular vibrational spectra.61, 80 Consequently, we adopted harmonic bonds and angles for better description of the dynamical behavior of the water molecules. The equilibrium bond length $r^{0}_{\text{OH}}$ and angle $\theta^{0}$ were tuned to ensure that the average values in the liquid phase align with experimental measurements at 298.15 K and 1 atm. The coefficients of bond length ($k_{b}$) and angle ($k_{a}$) were adjusted to produce accurate positions of the bond stretching and bending peaks in the IR spectrum.
@@
The electrostatic energy is calculated using ChargeNN charges. To facilitate simulations under periodic boundary conditions (PBC), Ewald summation is employed, utilizing the minimum image convention to construct the local environment for charge generation. It is important to note that the total charge of a cluster or unit cell could be non-zero due to the accumulative error of atomic charges. To ensure a zero net charge of the whole system, all atomic charges must be subtracted by a small constant ($\Delta q=\sum_{i}q_{i}/N$, where $N$ is the number of atoms) of approximately $\pm$0.0001 e. Additionally, non-homogeneous correction can be accomplished by considering the weights of atomic charges34 or integrating an explicit penalizing term to enforce electroneutrality into the loss function. There have been attempts to include charge transfer interactions with empirical charge transfer.136, 137, 138, 139 For our model that uses CT obtained quantum mechanically, the charge transfer term formulates a smooth ratio function of squared CT. The formula offers several merits, for instance, it is symmetric and the CT energy converges to zero in the absence of CT, preventing abrupt changes in potential energy around the CT cutoff. Additionally, the formula effectively captures the rapid intensification of CT interactions when initial charges are small, as well as the gradual attenuation due to the electron migration from equilibrium. Moreover, since the inter-water CT occurs away from the nuclear sites, the CT term naturally accounts for out-of-plane polarization, resulting in more accurate water conformations.
@@
As the inclusion of the CT term strengthens the short-range interactions, the traditional 12-6 LJ potential for Van der Waals (VDW) forces yielding a steep repulsive wall constrains inter-water distances to a compact range, which leads to an excessively high first peak in the oxygen-oxygen radial distribution function (RDF). Instead, we employ a 9-6 functional form with minor revisions to the original LJ formula, retaining the 12-6 LJ parameters $\varepsilon$ and $\sigma$ for compatibility with the existing force fields. In the process of the parameter tuning, the VDW equation was written as $U_{ij}^{\mathrm{vdw}}=A_{ij}/r^{9}_{ij}-B_{ij}/r^{6}_{ij}$ for convenience. The intermolecular parameters $k_{c}$, $k_{d}$, $A_{ij}$ and $B_{ij}$ were optimized for the accurate heat of vaporization, density and radial distribution functions (RDF) for liquid water at the ambient conditions. All parameters can be found in Table S1.
@@ 2.4 Implementation
Figure 2 illustrates the implementation of the ChargeNN water model for molecular dynamics (MD) simulations. In each MD step, atomic coordinates adjusted by the MD solver are utilized to compute the ML features interaction-classified functions $\mathbf{F}$ and their analytical derivatives with respect to nuclear positions $\mathrm{d}\mathbf{F}/\mathrm{d}\mathbf{R}$. The ML features are then fed into the pre-trained neural networks for the predictions of atomic charges and charge transfer. By employing auto-differentiation through back-propagation, the gradients of the charges concerning the features $\mathrm{d}q/\mathrm{d}\mathbf{F}$ can be conveniently evaluated. Subsequently, the predicted atomic charges and charge transfer are used to determine the potential energy as described in eq. 9, where total gradients are obtained using the chain rule:
@@
The ChargeNN program was written in Python with double precision, integrated with Tensorflow for handling neural networks. The algorithm has been parallelized based on Message Passing Interface (MPI) standard of version 3. The shared memory windows built in MPI-3’s remote memory access module enable significantly lower-latency data communications compared to traditional point-to-point communications.
@@ 3.1 Performance of the charge model
The ChargeNN model demonstrates exceptional data efficiency, requiring only 8000 training (H2O)25 clusters to achieve converged mean absolute errors (MAEs) for the remaining testing clusters, as shown in Figure S4. According to Figure 3, when trained on 8000 (H2O)25, the charge model yields testing MAEs of 0.0025 e for hydrogen and 0.0039 e for oxygen within the same system. These values are markedly lower than the respective MAEs of 0.0036 e and 0.0045 e for (H2O)23 predicted by a recent KRR water charge model trained with the same number of data points35. In addition, our model shows exceptionally high prediction reliability with the coefficients of determination ($R^{2}$) of 0.96 for hydrogen and 0.93 for oxygen, outperforming the 0.88 reported for the KRR model35. Notably, Figure 3 shows that the ChargeNN model considerably improves the transferability, only with a slow deterioration in accuracy as the water cluster size increases. For example, the models trained with (H2O)25 make accurate prediction of the charges for much larger (H2O)190, yielding MAEs of 0.0028 e for hydrogen and 0.0042 e for oxygen, both with $R^{2}$ values above 0.9. Figure S5 shows that predicting molecular charges by summing the ChargeNN atomic charges within each molecule is less satisfactory, with $R^{2}>0.92$ and marginally larger MAEs of $\sim$0.007 e due to the accumulative errors. This error can be improved in future by imposing constraints on the loss function to penalize the predicted molecular charges. For assessing ChargeNN’s performance in predicting charge transfer, we prepared charge transfer of 40000 water pairs extracted from the MD trajectory of (H2O)25. The ChargeNN trained on 20000 water pairs achieves an MAE of only 0.001 e and $R^{2}$ of 1.00 for the remaining testing pairs.
@@ 3.2 Performance of the water model
aThe origin was set to be the center of the nuclear charge of each molecule. bMolecules were wrapped to the unit cell. cThe gas-phase properties, interaction energy and melting temperature for SPC/FW were computed in this work, while other data were from ref.61. dSWM4-NDP results from ref.79, 80. eExperimental results from ref.140, 141, 142, 143, 144, 145.
@@
To assess the performance of the ChargeNN water model, we calculated various properties of water in gas, liquid, and solid states. Computational details are provided in the SI. We compare our results with those of SPC/FW61, which is the fixed charge counterpart of the 3-site ChargeNN model, and those with a widely used polarizable model SWM4-NDP79. As shown in Table 1, the ChargeNN’s predictions of water monomer and dimer properties agree well with experimental results, demonstrating its capability to accurately simulate gas-phase water clusters. Figure 4a illustrates that the angle $\theta_{A}$ of the water dimer obtained with ChargeNN (65) closely matches the experimental value of 58, in contrast to the unreasonable 22 by SPC/FW, highlighting the significant impact of polarization and charge transfer on the equilibrium geometries. Despite of the inclusion of intermolecular charge-transfer interactions in our model, enhanced electrostatics using high-order polarization may further improve the structure predictions.146 For instance, the O-O distance of the water dimer optimized by the multipolar AMOEBA (2.91 Å)67 aligns more closely with experimental measurement (2.98 Å) compared to ChargeNN (2.81 Å).
@@
Table 1 also demonstrates that the ChargeNN water model accurately reproduces bulk water properties at room temperature and standard pressure, such as average bond length and angle, density, interaction energy, vaporization heat, and diffusion constants. As shown in Figure 4b, ChargeNN produces an RDF for liquid oxygen-oxygen distances in good agreement with the MP2147 and early experimental data148. Nevertheless, the latest x-ray diffraction experiment149 has measured a less pronounced first peak of approximately 2.57 than the predicted one of 2.98, indicating the need for further parameter refinement to mitigate the “over-structured” hydrogen-bond network. The predicted infrared spectra in Figure 4c were obtained through a Fourier transformation of the autocorrelation function of the total dipole moments, in which the one produced by ChargeNN matches the experimental spectrum closely in both band positions and intensities. Notably, ChargeNN successfully captures the hydrogen-bond stretching peak at $\sim$200 cm-1, which is absent in the spectra by SPC/FW and SWM4-NDP. Our results and a previous study35 indicate that accurately describing hydrogen-bond stretching requires precise polarization and intermolecular CT. Additionally, ChargeNN retrieves the bending-libration combination band at $\sim$2150 cm-1, which is missing in the SPC/FW spectrum, further underscoring the importance of explicit polarization in predicting vibrational properties. Nevertheless, we observe that the O-H stretching band at around 3400 cm-1 is considerably narrower than the reference, which may be improved by replacing the harmonic intramolecular forces with anharmonic forces66 or reactive force fields150, 151.
@@
Figure 4: Assessment of the ChargeNN water model. (a) Geometries of water dimers optimized by SPC/FW and ChargeNN, compared to the experimental structure141. (b) The liquid radial distribution functions of the oxygen-oxygen distance obtained with ChargeNN, MP2147 and experiments 148, 149 under ambient conditions. (c) Liquid infrared spectra generated by SPC/FW, SWM4-NDP35, ChargeNN and experiment152 under ambient conditions. (d) Temperature dependent densities from SPC/FW153, SWM4-NDP81, ChargeNN and experiment154 at 1 atm.
@@
We computed the molecular dipole moments by setting the origin as the nuclear charge center of each partially charged water molecule, following the reported protocol132, 35. While ChargeNN accurately reproduces dipole moments for water monomer and dimer, as well as the liquid IR spectrum, it yields an average molecular dipole moment of approximately 1.76 Debye, which is notably lower than the experimental estimation of 2.9 Debye. While the choice of the origins may have a significant impact on the dipole moments of charged molecules, there could be parallel problems in the population analysis methods. A recent comprehensive study discovered that none of the selected charge models (Mulliken, Hirshfeld, QTAIM, RESP, ChelpG, Hirshfeld-I, and NPA) is capable of reproducing both total dipole moments and average molecular dipole moments of water clusters.132 Similarly, CM5 charges agree well with the reference values for total dipole moments, but fall in short to provide a reasonable average molecular dipole moment. Additionally, our water model underestimates the dielectric constant obtained from the total liquid dipole moments, which is largely due to the non-uniqueness of the dipole moment of an extended system according to the modern theory of polarization155.
@@
The accurate characterization across a range of temperatures is important to the performance of a water model. Figure 4d demonstrates that ChargeNN successfully predicts the temperature-dependent density of liquid water, with the temperature of maximum density around 280 K, very close to the experimental value of 277 K. In contrast, this prediction remains a challenge for the fixed charge model SPC/FW and other well-tuned 3-site models60, emphasizing the substantial improvement resulting from ChargeNN polarization. The polarizable SWM4-NDP also falls short in predicting the density variation with temperature. Additionally, Figure S8 presents that the computed temperature-dependence of vaporization enthalpy by ChargeNN agrees well with the reference. For the simulation of ice, we calculated the orientational tetrahedral order parameters of the equilibrium configurations across various temperatures, starting from a perfect hexagonal ($I_{h}$) ice structure. The ChargeNN yields a melting temperature of 288 K, close to the experimental value of 273 K. While this result is significantly better than SPC/FW at 190 K and SWM4-NDP at 185 K80, there is still room to fine-tune the parameters to achieve ice properties comparable to TIP4P/Ice57.
Figure 5: Timing performance of ChargeNN. (a) Elapsed time for force calculations on varying water box sizes using a single CPU core (AMD EPYC 7H12, 2.6GHz), comparing in-house ChargeNN and fixed point charges models with OpenMM’s SPC-FW and SWM4-NDP. (b) Speedup of ChargeNN calculations of periodic (H2O)10044 with respect to the number of the CPU cores.
As demonstrated in Figure 5a, ChargeNN exhibits exceptional computational efficiency, making it highly promising for large-scale water simulations. For instance, using a single CPU core for a PBC box containing 10044 water molecules, ChargeNN’s force calculation time is only one-eighth of that for the economical polarizable SWM4-NDP and merely five times that of the 3-site fixed charge model SPC-FW, both implemented in OpenMM for CPU platforms. Moreover, while the incorporation of machine-learned charges involves additional steps such as feature preparation, charge prediction, and charge derivatives, these ChargeNN-related steps only account for approximately 40% of the total time. Furthermore, the non-ChargeNN components (equivalent to fixed charge computations) are three times slower than OpenMM’s SPC-FW, indicating room for algorithmic improvement. In addition, Figure 5b reveals suboptimal parallel efficiency of the ChargeNN program, necessitating further optimization. We intend to transfer the algorithm to C++ for enhanced performance, with plans to adapt it to CUDA or OpenCL for extensive parallelization on graphics processing units (GPUs) in future implementations.